OpenStudy (anonymous):

What is eh equation of the line that passes through (-3,1) and has a slope of 3/5 but it needs to be in slope intercept form

OpenStudy (amistre64):

use point slope form and solve for y y - yo = m(x-xo) y = mx + (yo - mxo)